OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

y = - 1/7x +12/7

OpenStudy (anonymous):

Standard form: y = m*x + b where m is your slope and b is your y-intercept. m = Slope = rise/run = (y2 - y1) / (x2 - x1) = - 1/7 now we have y = (-1/7)x + b so input one of the points you have, lets use (-2,2) 2 = (-1/7)(-2) + b 2 = (2/7) + b b = 2 - (2/7) = (14/7) - (2/7) = 12/7 so then we put it all together to get y = -(1/7)x + 12/7 Sorry since I know Ishaan gave an answer already but simply giving answers doesn't help at all if you don't know how to use it/where it came from.